OpenStudy (anonymous):

Can someone help me please? Solve triangle ABC given that A = 56 degrees, B = 57 degrees, and b = 9. C=64 degrees, a=10.2, c=9.8 C=67 degrees, a=10.2, c=9.88 C=67 degrees, a=8.9, c=9.88 C=66 degrees, a=8.7, c=9.88

OpenStudy (anonymous):

 to find a and c, it's either you use the sine law or the cosine law

OpenStudy (ranga):

C = 180 - (A + B) A = 56, B = 57. Substitute above and solve for C Use the sine law to find a and c. a/sin(A) = b/sin(B) = c/sin(C) All angles are known and b = 9. So you can solve for a and c.